what is the slope of the line x=5
Found 2 solutions by palanisamy, CharlesG2:
Answer by palanisamy(496)

The equation of the given line is x = 5
This line is parallel to y-axis,
So its slope is infinity

Answer by CharlesG2(834)

what is the slope of the line x=5
that is a vertical line
vertical lines have undefined slopes
to demonstrate: m=slope=(y2-y1)/(x2-x1) with (x1,y1)=(5,5)
(x2,y2)=(5,0)
m=(0-5)/(5-5)=-5/0, you can not divide by 0, it is an infinite result, or in other words an undefined result